The combined area of two squares is 104 square meters. The sides of the larger square are five times as long as the sides of the smaller square. Find the dimensions of each of the squares.
Answer by sachi(548)

let the sides of the smaller square=x so area=x^2
the sides of the larger square=5x so the area=(5x)^2=25x^2
The combined area of two squares is 104 square meters.
so 25x^2+x^2=104
or 26x^2=104
or x^2=104/26=4
or x=2
the sides of the larger area=10
ans